# CoFrGeNet: Continued Fraction Architectures for Language Generation

### v1

@@ 1 Introduction
1) We propose novel continued fraction architectures for (causal) attention and FFNs as depicted in Figure 1. We call our architecture with both components replaced as Continued Fraction Generative Network (CoFrGeNet). We report results replacing either FFN or attention or both offering the possibility to the user of replacing only one or both of the components for their application. Even replacing one component can offer significant parameter and training time savings as seen in our experiments. 2) We propose an alternative representation for the ladders and derive custom formulas for the gradients that reduces the number of divisions from $d$ to a constant of just $1$ for a $d$-depth ladder. This greatly enhances both training and inference efficiency. 3) We propose a custom training schedule to update CoFrGeNet parameters. This is described in section 5. 4) We pre-train our models on OpenWebText (OWT) (Gokaslan et al., 2019), GneissWeb (Gohari et al., 2025) and the docling data mix (Team, 2024) showing that our models are competitive or outperform the corresponding Transformer models. We compare with Transformers since we are replacing its components making it a fair comparison. For an apples-to-apples comparison with other model architectures such as Mamba (Gu and Dao, 2024) one would want to replace its hidden state function with novel (to be designed) CoFrNet components, which would be a significant independent contribution in itself that we leave for future work.
@@ 2 Preliminaries
To construct networks out of continued fractions, we let the partial denominators $a_{k}$ be affine functions of an input $x$, $a_{k}=w_{k}x$, where $w_{k}$ is a row vector and a $1$ is prepended to the elements of $x$ so that the corresponding coefficient $w_{k0}$ is the intercept or “bias” term. We will often refer to a continued fraction with $a_{k}=w_{k}x$ as a (CoFrNet) “ladder”, and we will also construct ensembles of such ladders. Throughout the paper we denote the input or embedding dimension by $p$, the number of ladders in an ensemble by $L$, and sequence length by $l$, unless specified otherwise.
@@ 4.1.1 Replacement for Attention
Now considering the right architecture with two ladders (i.e. $L=2$) of depth $2$, a $L\times l$ (full) parameter matrix $F$ and Csoftmax to denote softmax applied causally (i.e. $i^{\text{th}}$ token is a convex combination of the first i1i-1 tokens) with notation from above we have attention weights given by,
@@
Figure 3: CoFrNet architecture to simulate FFNs – Cffn – in a transformer block. We create a gated non-expanded (i.e. $\alpha=1$) representation that we pass to the CoFrNet ladders. Here no transpose is taken and hence feature mixing in either direction does not interfere with causal generation which is why we have a linear layer on top. Again the collapsed implementation is described in section 4.2.
@@ 4.2 Architecture for Continued Fraction Ensembles and Continuant-Based Implementation
Architecture Let us denote by $y\in\mathbb{R}^{q}$ the output of a linear combination of $L$ ladders, where in general $q$ could be different from the input dimension $p$. We use a superscript $j$ to denote the partial denominators $a^{(j)}_{0},\dots,a^{(j)}_{d}$ corresponding to the $j$th ladder, where $a^{(j)}_{k}=w^{(j)}_{k}x$. Then based on (2), the $i$th output component $y_{i}$ is given by
@@
Advantages Using continuants to compute each continued fraction $\tilde{f}(a^{(j)})$ (6) and its gradient (9) requires only one division, by the same quantity $K_{d}$. As noted above, the reciprocal $1/K_{d}$ can be computed once and then reused in all ratios of continuants that are required. As seen from (5), all continuants up to $K_{d}$ can be computed recursively through $O(d)$ multiplications and additions. This continuants approach yields a major improvement in efficiency over the “literal” approach taken in the original CoFrNet work (Puri et al., 2021), which performs one division per layer following the standard representation of a continued fraction (1). The reduction from $d$ divisions to $1$ is especially significant when ladders are made deep. It applies to both inference and training, since backpropagation through a standard PyTorch implementation of (1) also requires $d$ divisions. It is widely known that divisions are significantly more expensive in current hardware — typically an order of magnitude slower — than multiplications or additions. Moreover, having to divide just once can result in better numerical stability.

### v2

@@ 1 Introduction
1) We propose novel continued fraction architectures for (causal) attention and FFNs as depicted in Figure 1. We call our architecture with both components replaced as Continued Fraction Generative Network (CoFrGeNet). We report results replacing either FFN or attention or both offering the possibility to the user of replacing only one or both of the components for their application. Even replacing one component can offer significant parameter and training time savings as seen in our experiments. 2) We propose an alternative representation for the ladders and derive custom formulas for the gradients that reduces the number of divisions from $d$ to a constant of just 11 for a $d$-depth ladder. This greatly enhances both training and inference efficiency. 3) We propose a custom training schedule to update CoFrGeNet parameters. This is described in section 5. 4) We pre-train our models on OpenWebText (OWT) (Gokaslan et al., 2019), GneissWeb (Gohari et al., 2025) and the docling data mix (Team, 2024) showing that our models are competitive or outperform the corresponding Transformer models. We compare with Transformers since we are replacing its components making it a fair comparison. For an apples-to-apples comparison with other model architectures such as Mamba (Gu and Dao, 2024) one would want to replace its hidden state function with novel (to be designed) CoFrNet components, which would be a significant independent contribution in itself that we leave for future work.
@@ 2 Preliminaries
To construct networks out of continued fractions, we let the partial denominators $a_{k}$ be affine functions of an input $x$, $a_{k}=w_{k}x$, where $w_{k}$ is a row vector and a 11 is prepended to the elements of $x$ so that the corresponding coefficient $w_{k0}$ is the intercept or “bias” term. We will often refer to a continued fraction with $a_{k}=w_{k}x$ as a (CoFrNet) “ladder”, and we will also construct ensembles of such ladders. Throughout the paper we denote the input or embedding dimension by $p$, the number of ladders in an ensemble by $L$, and sequence length by $l$, unless specified otherwise.
@@ 4.1.1 Replacement for Attention
Now considering the right architecture with two ladders (i.e. $L=2$) of depth 22, a $L\times l$ (full) parameter matrix $F$ and Csoftmax to denote softmax applied causally (i.e. $i^{\text{th}}$ token is a convex combination of the first i1i-1 tokens) with notation from above we have attention weights given by,
@@
Figure 3: CoFrNet architecture simulating FFNs – Cffn – in a transformer block. We create a gated non-expanded (i.e. $\alpha=1$) representation that we pass to the CoFrNet ladders. No transpose is taken and hence feature mixing in either direction does not interfere with causal generation which is why we have a linear layer on top. Again the collapsed implementation is described in section 4.2.
@@ 4.2 Architecture for Continued Fraction Ensembles and Continuant-Based Implementation
Architecture Let us denote by $y\in\mathbb{R}^{q}$ the output of a linear combination of $L$ ladders, where in general $q$ could be different from the input dimension pp. We use a superscript $j$ to denote the partial denominators $a^{(j)}_{0},\dots,a^{(j)}_{d}$ corresponding to the $j$th ladder, where $a^{(j)}_{k}=w^{(j)}_{k}x$. Then based on (2), the $i$th output component $y_{i}$ is given by
@@
Advantages Using continuants to compute each continued fraction $\tilde{f}(a^{(j)})$ (6) and its gradient (9) requires only one division, by the same quantity $K_{d}$. As noted above, the reciprocal $1/K_{d}$ can be computed once and then reused in all ratios of continuants that are required. As seen from (5), all continuants up to $K_{d}$ can be computed recursively through $O(d)$ multiplications and additions. This continuants approach yields a major improvement in efficiency over the “literal” approach taken in the original CoFrNet work (Puri et al., 2021), which performs one division per layer following the standard representation of a continued fraction (1). The reduction from $d$ divisions to 11 is especially significant when ladders are made deep. It applies to both inference and training, since backpropagation through a standard PyTorch implementation of (1) also requires $d$ divisions. It is widely known that divisions are significantly more expensive in current hardware — typically an order of magnitude slower — than multiplications or additions. Moreover, having to divide just once can result in better numerical stability.